change from a mixed number to an improper fraction 12 8/9
Answer by jim_thompson5910(35256)

12 8/9 = 12+8/9

12 8/9 = 12(9/9)+8/9

12 8/9 = 108/9+8/9

12 8/9 = (108+8)/9

12 8/9 = 116/9

So the answer is 116/9

--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------